

********************************************************************
UARS WEEKLY STATUS REPORT - 19 AUG 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM's final data was gathered on August 1. No more operations are
planned. We are preparing our (perhaps) final data version.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005099 has been processed and is waiting for DAAC archival.
HEPSB data files are version 1. HEPSB data files between 1991274 and
1991366 have been processed and are waiting for DAAC archival.
Processing of all PEM DAAC data files has been suspended until the
DAAC catches up with archiving all processed data files.
AXIS1 data file formats are beginning design.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 06 August 2005 (Orbit 76099, GMT Day 218) through
12 August 2005 (Orbit 76204, GMT Day 224).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
NONE THIS REPORT PERIOD.
A Reverse-to-Forward Yaw Around maneuver (number 142) was successfully
executed in Orbit 76155 beginning 221/ 18:11:45Z and completing at
221/ 18:48:55Z on 09 August 2005. All instruments were in their off
or standby mode (except for SOLSTICE in day-only). Due to the SA
position at 265 degrees, the effective night length of the spacecraft
after the maneuver was approximately 53.52 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.40 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 21.55 and 14.45 percent
respectively. The next (forward-to-reverse) YAW is scheduled for 17
September 2005 (DOY 260).
The beta angle ranged between the angles of 15.4 degrees to 10.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 66 degrees on 29 August 2005 (DOY 242).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 6.89 to 6.83 degrees C. Additional
battery performance data for each battery on 05 August 2005 and 12
August 2005 are provided in Appendix D.
The clock error ranged from 10.1 to -10.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 07
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT used the VT7 charge level for 2 orbits during the yaw. This
week the FOT continued to closely monitor the batteries. The
differential voltages on battery 2 have dropped a bit from last week.
It now ranges between -240mV and -350mV this week. The max
temperature on battery 2 has come down a bit since the yaw.
MEETINGS
None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 September 2005 (DOY 260).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 August 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
76143 220/2227 05-250 PSU DISABLE TMON #7 (BAT 2 DIFV)
76151 221/1103 05-251 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76154 221/1538 05-252 MPS CHANGE VT LEVEL TO 7
76156 221/1918 05-253 OBC YAW
76516 221/1949 05-254 MPS CHANGE VT LEVEL TO 6
76159 222/0030 05-255 OBC ATC NOOPS (DELETED PASSES)
76172 222/2131 05-256 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 August 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 August 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
76151 221/1127 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76172 222/2132 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 August 2005
05 August 2005 - GMT Day 217 Beta = 15.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.42/- 3.68 +00.0/+00.0 04.8 0.00 0.0
2 -112.0/-425.6 +12.10/+11.07 +32.8/-17.6 25.92 1.08 21.7
3 44.8/-700.0 +4.50/+ 3.87 +19.2/-10.4 25.76 1.05 12.7
12 August 2005 - GMT Day 224 Beta = 10.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -3.05/ -4.31 +00.0/+00.0 04.64 0.00 0.0
2 -89.6/-280.0 +11.07/+10.06 +31.6/-16.4 26.24 1.09 20.9
3 44.8/-700.0 +4.18/+ 3.23 +20.8/-09.6 26.08 1.05 13.8
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
76154 221/1538 LV6 -> LV7
76516 221/1949 LV7 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 August 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-039 FOT 76162 222/0506 40451 WS Chain Failover
05-040 FOT 76176 223/0244 40465 TLM HITS - WEATHER AT GUAM
05-041 FOT 76183 223/1443 40470 TLM DROPOUT - WEATHER AT GUAM
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-032 FOT See List July 05 30832 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-039 FOT 76162 222/0506 40451 WS Chain Failover
05-040 FOT 76176 223/0244 40465 TLM HITS - WEATHER AT GUAM
05-041 FOT 76183 223/1443 40470 TLM DROPOUT - WEATHER AT GUAM